# From tools to thieves: Measuring and understanding public perceptions of AI through crowdsourced metaphors

### v1

@@ Abstract.
How has the public responded to the increasing prevalence of artificial intelligence (AI)-based technologies? We investigate public perceptions of AI by collecting over 12,000 responses over 12 months from a nationally representative U.S. sample. Participants provided open-ended metaphors reflecting their mental models of AI, a methodology that overcomes the limitations of traditional self-reported measures. Using a mixed-methods approach combining quantitative clustering and qualitative coding, we identify 20 dominant metaphors shaping public understanding of AI. To analyze these metaphors systematically, we present a scalable framework integrating language modeling (LM)-based techniques to measure key dimensions of public perception: anthropomorphism (attribution of human-like qualities), warmth, and competence. We find that Americans generally view AI as warm and competent, and that over the past year, perceptions of AI’s human-likeness and warmth have significantly increased ($+34\%,r=0.80,p<0.01;+41\%,r=0.62,p<0.05$). Furthermore, these implicit perceptions, along with the identified dominant metaphors, strongly predict trust in and willingness to adopt AI ($r^{2}=0.21,0.18,p<0.001$). We further explore how differences in metaphors and implicit perceptions—such as the higher propensity of women, older individuals, and people of color to anthropomorphize AI—shed light on demographic disparities in trust and adoption. In addition to our dataset and framework for tracking evolving public attitudes, we provide actionable insights on using metaphors for inclusive and responsible AI development.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYcopyright: noneisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Perceptions about new technologies, however, can be difficult to capture using traditional survey measures (Gallagher and Updegraff, 2012; Brugman et al., 2017; Jiang et al., 2021; Jensen et al., 2024). People often struggle to verbalize their nuanced perceptions of complex sociotechnical systems like AI, particularly when they have less experience with them (Lee et al., 2021). To overcome these limitations, we study public perceptions of AI by analyzing the metaphors they use to describe AI (Demir and Güraksın, 2022). People have long used metaphors to communicate complex ideas, like the turbulence of emotions (“rollercoaster of feelings”) or the value of digital resources (“data is the new oil”) (Lakoff and Johnson, 1980). We conducted an online study where participants were asked to provide their metaphors of AI, collecting responses from a nationally representative sample of over 12,000 Americans from May 2023 (shortly after the mainstream adoption of ChatGPT (Wu et al., 2023)) to August 2024. The collection of these metaphor data over this critical period provides insights into how the public interprets and understands the rise of generative AI and public-facing tools like ChatGPT.
First, we identified the dominant metaphors people used throughout this time period to describe AI (e.g., “AI is a tool”, “AI is a thief”) using a combination of automatic clustering and qualitative refinement. Beyond the explicit descriptions of the metaphors, we also demonstrate how these metaphors provide a window into people’s implicit perceptions of AI. We measure implicit perceptions from the metaphors in two ways: 1) the extent to which participants anthropomorphized AI, using a method measuring implicit framings from language model (LM)-based probabilities (Cheng et al., 2024), and 2) how warm and competent they perceive AI, using semantic axes constructed with LM-based embeddings (Fraser et al., 2021). These constructs are important predictors of people’s experiences with AI: Anthropomorphism, or the attribution of human-like qualities to AI, has come to the fore of discussions around societal impacts of AI (Salles et al., 2020; Shanahan et al., 2023; Troshani et al., 2021; Cheng et al., 2025) as scholars have highlighted the consequences of users becoming overly dependent on human-like AI (Abercrombie et al., 2023; Chiesurin et al., 2023; Chan et al., 2023; Inie et al., 2024) and failing to consider the implications of disclosing sensitive information (Ischen et al., 2020; Citron and Solove, 2022). Warmth and competence are integral dimensions of how people form impressions of social actors (Fiske et al., 2007; Mieczkowski et al., 2019). The extent to which people believe that others have warm, positive intentions and are capable of carrying out given tasks are strong predictors of people’s attitudes, such as their trust in new technologies and willingness to adopt them in the future (McKee et al., 2023; Li et al., 2024; Khadpe et al., 2020; Zhou et al., 2024).
@@
Our study provides a fine-grained understanding of how public perceptions of AI have emerged over time and across individuals. We collect a dataset of over 12,000 metaphors of AI over 12 months and use mixed methods to identify 20 dominant metaphors. We provide a scalable framework for measuring implicit perceptions from the metaphors. We use this framework to demonstrate how temporal shifts in metaphors and implicit perceptions correspond to changes in people’s attitudes, and explore how demographic differences among these reveal harmful consequences of AI deployment. Our framework is quantitative and scalable, and thus can be easily adapted to new data to capture evolutions in public perception (Fast and Horvitz, 2017; Zhang and Dafoe, 2019).  [^2] Our code is available at https://anonymous.4open.science/r/ai_metaphors-8008, and the full dataset is available upon request.
@@ 2.2. Why metaphors matter
Because they help people make sense of unfamiliar concepts, metaphors have long been used to help people understand new technologies. Early metaphors of the Internet as a “superhighway” that could connect users to diverse digital destinations (Fluckiger, 1996) and the later metaphor of “surfing the web” that suggests the Internet as a vehicle for exploration (Maglio and Matlock, 1998). Similarly, understanding the metaphors that people use to describe AI may reveal how the public is responding to the proliferation of AI-based technologies, and how they understand the potential values and uses that AI technologies may afford. Already, clear differences are emerging in how people think about AI. For instance, some people describe AI is a powerful tool that should be integrated into the workforce (Hazarika, 2020; Eubanks, 2022; Stade et al., 2024). Others view AI as an assistant or companion that can be trusted to process personal and professional challenges (Danaher, 2018). However, many people view AI to be a threat to their livelihoods, invoking fears of replacement and may therefore instead advocate for restrictions on AI (Kochhar, 2023). The metaphors people use to describe AI provide insight into how people feel about using it in their own lives, in the context of their work, and in society at large.
@@ 2.3. Implicit perceptions of AI
People’s implicit perceptions play a powerful role in shaping human interactions with technology. Even when people interact with the same system, they can interpret the qualities of the system in vastly different ways that affect their trust and engagement. Understanding these perceptual differences is vital for building and deploying AI responsibly. Because the inner workings of these systems are often “black boxes”, individuals rely on their intuitive beliefs when communicating with and through AI (Ytre-Arne and Moe, 2021). Similarly, the effects of interacting with AI depend on how people perceive and interpret their experiences with these complex socio-technical systems. Previous research on generative AI has emphasized the importance of considering several dimensions of these perceptions, which we measure using our quantitative framework:
1) Anthropomorphism: Anthropomorphism, or the attribution of human-like characteristics in a non-human entity, spans a broad range of qualities, including cognitive, behavioral, and emotional traits (Epley et al., 2007; DeVrio et al., 2025). In the context of AI, this can manifest in applications like conversational assistants (Gabriel et al., 2024), relationship-building chatbots (Brandtzaeg et al., 2022), and AI-generated simulations of individuals (McIlroy-Young et al., 2022). Anthropomorphism of AI is contentious at the individual and societal levels, as it raises concerns around assigning moral responsibility to AI (Friedman and Kahn Jr, 2007), overreliance on AI (Watson, 2019; Zarouali et al., 2021), and degrading emotional and social relations (Akbulut et al., 2024; Maeda and Quan-Haase, 2024; Shteynberg et al., 2024; Laestadius et al., 2022).
@@ 2.4. Attitudes toward AI: Trust and adoption
We focus on two core aspects of public attitudes toward AI: trust in AI and willingness to adopt AI. First, these variables are important because they are necessary for effective and productive integration of AI into society (Afroogh et al., 2024). Lack of trust and unwillingness to adopt AI may lead to under-use and missed opportunities to improve people’s lives(Choudhury and Shamszare, 2023). Moreover, it may result in public fear or backlash (Lukyanenko et al., 2022) that diverts attention toward unfounded, hypothetical risks while neglecting the real and ongoing harms that AI technologies already impose on various populations (Chien and Danks, 2024; Tiku, 2022; Hunter, 2023). At the same time, it is also critical for users to not over-trust AI and understand its biases and limitations (Weidinger et al., 2022), as failing to do so can lead to harms related to over-reliance (Chiesurin et al., 2023; Abercrombie et al., 2023).
@@ 3.1. Participants and Recruitment
We recruited 12,933 participants from the crowdsourcing platform Prolific between May 2023 and August 2024, with approximately 1,000 individuals recruited each month, as part of a larger project understanding Americans’ experiences with AI. We note that over the 16-month period, we were unable to collect data during July 2023 and May - July 2024 due to technical issues; thus, we have responses over 12 months total. All survey procedures were approved by the BetterUp Institutional Review Board. We assessed the degree to which the samples each month were nationally representative of the United States population by using the American National Election Studies’ raking algorithms (Baker et al., 2013; Pasek, 2016). The analysis revealed that our data were nationally representative of the US population with respect to gender, ethnicity, education, and age overall, as none of the variables differed by a margin of more than .5%. For full participant demographics and month-by-month splits, please see Appendix A.
@@ 3.2. Procedure and Measures
We elicited metaphors from participants by asking them: “Some people use metaphors to describe abstract concepts, like AI. What is the best metaphor for how AI works?” They were encouraged to use their own words, and were assured that there were no wrong answers. We specified that the task was not about assessing their comprehension of the details of how AI works, but rather about understanding their thoughts (“This is not about accuracy as much as understanding how you are thinking”). Participants entered their answer into an open text box in the form “AI is like __________because___________.” Next, participants completed a series of survey measures related to their experiences with and attitudes toward AI. We assessed the frequency of each individual’s AI use by asking them to indicate if they had heard of or used 8 of the most commonly used, consumer-facing tools: ChatGPT, DALL-E, Claude, Bard, Anthropic, Midjourney, Gemini, and Perplexity. They could also write-in the names of additional AI tools that they used. Trust in AI was measured using a 3-item survey measure adapted from past work investigating trust in human vs. AI-generated content (Jakesch et al., 2019), building on the Organizational Model of Trust which is widely used to study trust in technologies and products (Mayer, 1995). Finally, we assessed participants willingness to adopt AI with a 5-item survey measure of their behavioral intentions to engage with AI in the future. Full details of these survey items are in Appendix B.
@@ 4.1. Topic modeling to identify dominant metaphors
To identify thematic clusters in the metaphors that participants provide, we use a mixed-methods approach: 1) automatic clustering using embeddings, 2) manual refinement of clusters, and 3) outlier analysis.
1) Automatic clustering: First, we removed frequently occurring words (e.g., “AI” and “artificial intelligence”) from all metaphors to ensure that the identified clusters were based on meaning rather than differences in phrasing. We then use the sentence embedding model all-mpnet-base-v2 (Reimers and Gurevych, 2019) to generate a 768-dimensional embedding $e({m_{p}})$ that captures the meaning of each participant $p$’s metaphor of AI $m_{p}$. Following standard practice (Kirk et al., 2024), these representations were reduced to 20 dimensions using UMAP (McInnes et al., 2018) and normalized using the L2 norm to reduce complexity prior to clustering. We then clustered these representations with HDBScan, a clustering algorithm that groups similar metaphors without forcing every metaphor into a group or pre-defining a number of groups (McInnes et al., 2017). The clustering was performed using the BERTopic Python package (Grootendorst, 2022). This process categorized about 80% of the metaphors in one of 50 clusters, while the remaining 20% were considered outliers.
@@ 4.2.1. Measuring anthropomorphism
(1) $A(s_{x})=\log\frac{P_{\textsc{human}}(s_{x})}{P_{\textsc{non-human}}(s_{x})}\text{where }P_{\textsc{human}}(s_{x})=\sum_{w\in\text{he, she}}P(w),P_{\textsc{non-human}}(s_{x})=\sum_{w\in\text{it}}P(w),$
where $P(w)$ is the model’s outputted probability of replacing the mask with the word $w$.
@@
where $w_{i}/w^{\prime}_{j}$ is a word in the set of words positively/negatively associated with that dimension respectively. We again use all-mpnet-base-v2 to compute the embeddings. For each metaphor $m_{p}$ embedded as $e(m_{p})$, we compute its warmth and competence as the cosine similarity of the embedding $e(m_{p})$ to this axis, i.e.,
(4) $\text{Warmth}(m_{p})=\cos(e(m_{p}),A_{\text{warmth}}),\text{Competence}(m_{p})=\cos(e(m_{p}),A_{\text{competence}}),$
resulting in two scores (each between -1 and 1) that represent the metaphor’s warmth and competence respectively.

### v2

@@ Abstract.
How has the public responded to the increasing prevalence of artificial intelligence (AI)-based technologies? We investigate public perceptions of AI by collecting over 12,000 responses over 12 months from a nationally representative U.S. sample. Participants provided open-ended metaphors reflecting their mental models of AI, a methodology that overcomes the limitations of traditional self-reported measures by capturing more nuance. Using a mixed-methods approach combining quantitative clustering and qualitative coding, we identify 20 dominant metaphors shaping public understanding of AI. To analyze these metaphors systematically, we present a scalable framework integrating language modeling (LM)-based techniques to measure key dimensions of public perception: anthropomorphism (attribution of human-like qualities), warmth, and competence. We find that Americans generally view AI as warm and competent, and that over the past year, perceptions of AI’s human-likeness and warmth have significantly increased ($+34\%,r=0.80,p<0.01;+41\%,r=0.62,p<0.05$). These implicit perceptions, along with the identified dominant metaphors, strongly predict trust in and willingness to adopt AI ($r^{2}=0.21,0.18,p<0.001$). Moreover, we uncover systematic demographic differences in metaphors and implicit perceptions, such as the higher propensity of women, older individuals, and people of color to anthropomorphize AI, which shed light on demographic disparities in trust and adoption. In addition to our dataset and framework for tracking evolving public attitudes, we provide actionable insights on using metaphors for inclusive and responsible AI development.
copyright: acmlicensedjournalyear: 2025conference: ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT ’25), June 23–26, 2025, Athens, Greececopyright: none
@@ 1. Introduction
Perceptions about new technologies, however, can be difficult to capture using traditional survey measures (Gallagher and Updegraff, 2012; Brugman et al., 2017; Jiang et al., 2021; Jensen et al., 2024). People often struggle to verbalize their nuanced perceptions of complex sociotechnical systems like AI, particularly when they have less experience with them (Lee et al., 2021). To overcome these limitations, we study public perceptions of AI by analyzing the metaphors they use to describe AI (Demir and Güraksın, 2022). People have long used metaphors to communicate complex ideas, like the turbulence of emotions (“rollercoaster of feelings”) or the value of digital resources (“data is the new oil”) (Lakoff and Johnson, 1980). We conducted an online study where participants were asked to provide their metaphors of AI, collecting responses from a nationally representative sample of over 12,000 Americans from May 2023 (shortly after the mainstream adoption of ChatGPT (Wu et al., 2023)) to August 2024. The collection of these metaphor data over this critical period provides insights into how the public interprets and understands the rise of generative AI and public-facing tools.
First, we identified the dominant metaphors people used throughout this time period to describe AI (e.g., “AI is a tool”, “AI is a thief”) using a combination of automatic clustering and qualitative refinement. Beyond the explicit descriptions of the metaphors, we also demonstrate how these metaphors provide a window into people’s implicit perceptions of AI. We measure implicit perceptions from the metaphors in two ways: 1) the extent to which participants anthropomorphized AI, using a method measuring implicit framings from language model (LM)-based probabilities (Cheng et al., 2024), and 2) how warm and competent they perceive AI, using semantic axes constructed with LM-based embeddings (Fraser et al., 2021). These constructs are important predictors of people’s experiences with AI: Anthropomorphism, or the attribution of human-like qualities to AI, has come to the fore of discussions around societal impacts of AI (Salles et al., 2020; Shanahan et al., 2023; Troshani et al., 2021; Cheng et al., 2025b) as scholars have highlighted the consequences of users becoming overly dependent on human-like AI (Abercrombie et al., 2023; Chiesurin et al., 2023; Chan et al., 2023; Inie et al., 2024) and failing to consider the implications of disclosing sensitive information (Ischen et al., 2020; Citron and Solove, 2022). Warmth and competence are integral dimensions of how people form impressions of social actors (Fiske et al., 2007; Mieczkowski et al., 2019). The extent to which people believe that others have warm, positive intentions and are capable of carrying out given tasks are strong predictors of people’s attitudes, such as their trust in new technologies and willingness to adopt them in the future (McKee et al., 2023; Li et al., 2024; Khadpe et al., 2020; Zhou et al., 2024).
@@
Our study provides a fine-grained understanding of how public perceptions of AI have emerged over time and across individuals. We collect a dataset of over 12,000 metaphors of AI over 12 months and use mixed methods to identify 20 dominant metaphors. We provide a scalable framework for measuring implicit perceptions from the metaphors. We use this framework to demonstrate how temporal shifts in metaphors and implicit perceptions correspond to changes in people’s attitudes, and explore how demographic differences among these reveal harmful consequences of AI deployment. Our framework [^2] Our code is available at https://github.com/myracheng/ai_metaphors, and our dataset is available at https://osf.io/jers8. is quantitative and scalable, and thus can be easily adapted to new data to capture evolutions in public perception (Fast and Horvitz, 2017; Zhang and Dafoe, 2019).
@@ 2.2. Why metaphors matter
Because they help people make sense of unfamiliar concepts, metaphors have long been used to help people understand new technologies. Early metaphors of the Internet as a “superhighway” that could connect users to diverse digital destinations (Fluckiger, 1996) and the later metaphor of “surfing the web” that suggests the Internet as a vehicle for exploration (Maglio and Matlock, 1998). Similarly, understanding the metaphors that people use to describe AI may reveal how the public is responding to the proliferation of AI-based technologies, and how they understand the potential values and uses that AI technologies may afford. Clear differences are emerging in how people think about AI. For instance, some people describe AI is a powerful tool that should be integrated into the workforce (Hazarika, 2020; Eubanks, 2022; Stade et al., 2024). Others view AI as an assistant or companion that can be trusted to process personal and professional challenges (Danaher, 2018). However, many people view AI to be a threat to their livelihoods, invoking fears of replacement and may therefore instead advocate for restrictions on AI (Kochhar, 2023). Previous work has also demonstrated the impact of human versus non-human metaphors in conversational agent design (Desai and Twidale, 2023; Jung et al., 2022). The metaphors people use to describe AI provide insight into how people feel about using it in their own lives, in the context of their work, and in society at large.
@@ 2.3. Implicit perceptions of AI
People’s implicit perceptions play a powerful role in shaping human interactions with technology. Even when people interact with the same system, they can interpret the qualities of the system in vastly different ways that affect their trust and engagement. Understanding these perceptual differences is vital for building and deploying AI responsibly. Because the inner workings of these systems are often “black boxes”, individuals rely on their intuitive beliefs when communicating with and through AI (Ytre-Arne and Moe, 2021). Similarly, the effects of interacting with AI depend on how people perceive and interpret their experiences with these complex socio-technical systems. Previous research on generative AI has emphasized the importance of considering key dimensions of perceptions, which we measure using our quantitative framework:
1) Anthropomorphism: Anthropomorphism, or the attribution of human-like characteristics in a non-human entity, spans a broad range of qualities, including cognitive, behavioral, and emotional traits (Epley et al., 2007; DeVrio et al., 2025; Cheng et al., 2025c, a). In the context of AI, this can manifest in applications like conversational assistants (Gabriel et al., 2024), relationship-building chatbots (Brandtzaeg et al., 2022), and AI-generated simulations of individuals (McIlroy-Young et al., 2022; Olteanu et al., 2025). Anthropomorphism of AI is contentious at the individual and societal levels, as it raises concerns around assigning moral responsibility to AI (Friedman and Kahn Jr, 2007), overreliance on AI (Watson, 2019; Zarouali et al., 2021), and degrading emotional and social relations (Akbulut et al., 2024; Maeda and Quan-Haase, 2024; Shteynberg et al., 2024; Laestadius et al., 2022).
@@ 2.4. Attitudes toward AI: Trust and adoption
We focus on two core aspects of public attitudes toward AI: trust in AI and willingness to adopt AI. A certain level of trust is necessary for effective and beneficial integration of AI into society (Afroogh et al., 2024); lack of trust and unwillingness to adopt AI may lead to under-use and missed opportunities to improve people’s lives (Choudhury and Shamszare, 2023). However, excessive trust or premature adoption can also be dangerous, leading to over-reliance, inaccurate expectations of AI capabilities, and other harms (Chiesurin et al., 2023; Abercrombie et al., 2023). Over-trust can also obscure the real, ongoing harms that AI systems already impose, including the amplification of social inequalities (Birhane, 2021; Bianchi et al., 2023), surveillance and erosion of privacy (Kalluri et al., 2023), and the marginalization of vulnerable groups (Weidinger et al., 2022; Chien and Danks, 2024; Tiku, 2022; Hunter, 2023). Moreover, public fear or backlash fueled by misplaced trust can divert attention away from addressing these concrete harms, instead focusing on speculative or exaggerated risks (Lukyanenko et al., 2022). Understanding these variables is thus critical to shaping societal outcomes of AI.
@@ 3.1. Participants and Recruitment
We recruited 12,933 participants from the crowdsourcing platform Prolific between May 2023 and August 2024, with approximately 1,000 individuals recruited each month, as part of a larger project understanding Americans’ experiences with AI. We note that over the 16-month period, we were unable to collect data during July 2023 and May - July 2024 due to technical issues; thus, we have responses over 12 months total. All survey procedures were approved by the BetterUp Institutional Review Board. We assessed the degree to which the samples each month were nationally representative of the United States population by using the American National Election Studies’ raking algorithms (Baker et al., 2013; Pasek, 2016). The analysis revealed that our data were nationally representative of the US population with respect to gender, ethnicity, education, and age overall, as none of the variables differed by a margin of more than .5% in any month. For full demographic details, please see Appendix A.
@@ 3.2. Procedure and Measures
We elicited metaphors from participants by asking them: “Some people use metaphors to describe abstract concepts, like AI. What is the best metaphor for how AI works?” They were encouraged to use their own words, and were assured that there were no wrong answers. We specified that the task was not about assessing their comprehension of the details of how AI works, but rather about understanding their thoughts (“This is not about accuracy as much as understanding how you are thinking”). We chose to give this prompt with relatively little context, rather than a specific hypothetical scenario, to capture a wide range of responses. Participants entered their answer into an open text box in the form “AI is like __________because___________.” Next, participants completed a series of survey measures related to their experiences with and attitudes toward AI. We assessed the frequency of each individual’s AI use by asking them to indicate if they had heard of or used 8 of the most commonly used, consumer-facing tools: ChatGPT, DALL-E, Claude, Bard, Anthropic, Midjourney, Gemini, and Perplexity. They could also write-in the names of additional AI tools that they used. Trust in AI was measured using a 3-item survey measure adapted from past work investigating trust in human vs. AI-generated content (Jakesch et al., 2019), building on the Organizational Model of Trust widely used to study trust in technologies and products (Mayer, 1995). Finally, we assessed participants willingness to adopt AI with a 5-item survey measure of their behavioral intentions to engage with AI in the future. Full details of these survey items are in Appendix B.
@@ 4.1. Topic modeling to identify dominant metaphors
To identify thematic clusters in participants’ metaphors, we use a mixed-methods approach: 1) automatic clustering using embeddings, 2) manual refinement of clusters, and 3) outlier analysis.
1) Automatic clustering: First, we removed frequently occurring words (e.g., “AI” and “artificial intelligence”) from all metaphors to ensure that the identified clusters were based on meaning rather than differences in phrasing. We then use the sentence embedding model all-mpnet-base-v2 (Reimers and Gurevych, 2019) to generate a 768-dimensional embedding $e({m_{p}})$ that captures the meaning of each participant $p$’s metaphor of AI $m_{p}$. Following prior work (Kirk et al., 2024) and based on qualitative examination of the clusters generated from varying dimensionalities, these representations were reduced to 20 dimensions using UMAP (McInnes et al., 2018) and normalized using the L2 norm to reduce complexity prior to clustering. We then clustered these representations with HDBScan, a clustering algorithm that groups similar metaphors without forcing every metaphor into a group or pre-defining a number of groups (McInnes et al., 2017). The clustering was performed using the BERTopic Python package (Grootendorst, 2022). This process categorized about 80% of the metaphors in one of 50 clusters, while the remaining 20% were considered outliers.
@@ 4.2.1. Measuring anthropomorphism
$A(s_{x})=\log\frac{P_{\textsc{human}}(s_{x})}{P_{\textsc{non-human}}(s_{x})}\text{, where}$
(1) $P_{\textsc{human}}(s_{x})=\sum_{w\in\text{he, she}}P(w),P_{\textsc{non-human}}(s_{x})=\sum_{w\in\text{it}}P(w),$
@@
where $w_{i}/w^{\prime}_{j}$ is a word in the set of words positively/negatively associated with that dimension respectively. We again use
all-mpnet-base-v2 to compute the embeddings. For each metaphor $m_{p}$ embedded as $e(m_{p})$, we compute its warmth and competence as the cosine similarity of the embedding $e(m_{p})$ to this axis, i.e.,
(4) $\begin{split}\text{Warmth}(m_{p})&=\cos(e(m_{p}),A_{\text{warmth}}),\\ \text{Competence}(m_{p})&=\cos(e(m_{p}),A_{\text{competence}})\end{split}$
resulting in two scores (each between -1 and 1) that represent the metaphor’s warmth and competence respectively. Examples of metaphors with different scores for warmth and competence are in Table 1.
Positive warmth, negative competence: “It grinds my gears” “Like a turtle trying to run like a rabbit.” “Its a computer predicting the next word It has no understanding” ”Shoddy thief” Positive warmth, positive competence: “AI is like having a ‘phone a friend’ lifeline on the TV gameshow, ‘Who wants to be a millionaire.’ You can ask for assistance from a more knowledgeble entity for personal assistance or gain.” “AI is like a wise old friend because it knows a lot and can give very helpful advice.” “It makes me think of a fairy god mother. All knowing and there to help you out” “It’s like a robot that can learn over time”
Negative warmth, negative competence: “It’s like a dog. On one hand, it can be gentle and loving. However, you don’t know if it may suddenly bite.” “Like an imaginary friend” “AI is like Your uncle is who right 50% of time. because They have a lot to say but don’t understand the context.” “AI is like having a forever friend because they cant get up and leave” Negative warmth, positive competence: “A detective that collects clues from different recourses, analyzes them and uses that information to make a prediction or solve a problem” “Its a computer system that is loaded with a bunch of data. ALl the data is tied to key words so when the key word us used a bunch of data with that tag is pulled up.” “It’s like a copy machine that takes info and tries to replicate it” “Putting human knowledge into a system and then having it spit out I got stood at you. It could be in the form or writing or even art.”
Table 1. Metaphors with varying levels of warmth and competence (participants’ responses, unedited). We find that the vast majority of metaphors have positive warmth and positive competence.